Other things being equal, how could the price of a constant growth stock be affected if there is a decrease in its rate of return, rS?

Group of answer choices

No change.

Lower.

Higher.
Answer #1

Option (c) is correct

Price will be higher.

As per Gordon model, share price is given by:

Share price = D1 / k -g

where, D1 is next years' dividend, k is the required rate of return  and g is the growth rate

Suppose ,

D1 = $100, K = 10% and g = 5%

Putting the values in the above formula, we get

Share price = $100 / 10% - 5%

Share price = $100 / 5% = $2000

Now, suppose rate of return i.e. k decreases to 6%,

Putting the values in the above formula, we get

Share price = $100 / 6% - 5%

Share price = $100 / 1% = $10000

So, when the rate of return decreases, share price will increase.